OpenStudy (anonymous):

Given: f(x) = 3 - x; g(x) = -2x Find g[f(x)]. 2x - 6 -x - 6 2x + 3

OpenStudy (anonymous):

this means that you must take g(x) and put 3-x for x in g(x)

OpenStudy (anonymous):

g(x)=-2x and f(x)=3-x so it would be g(f(x))=-2(3-x)

OpenStudy (anonymous):

ok Thanks I get it!!! 11:11 Make a wish!!!

OpenStudy (anonymous):

11:11 where?

OpenStudy (anonymous):

lol Its just something I do, when Ive been helped!!

OpenStudy (anonymous):

okay because I am in california and its 7:52pm

OpenStudy (anonymous):

LoL!!!